OpenStudy (anonymous):

The length of each side of a square wooden box, in inches, is represented by the expression 6a3. The volume of the box, in cubic inches, is (6a3)3. Which simplified expression represents the volume of the box?

OpenStudy (cwrw238):

(6a^3)^3 = 6^3* a^(3*3)

OpenStudy (cwrw238):

can you continue with this?

OpenStudy (anonymous):

i got 216a^27 ?

OpenStudy (cwrw238):

216 is is correct, but (a^3)^3 = a^(3*3) = a^9

OpenStudy (cwrw238):

one of rule of indices is( a^m)^n = a^mn

OpenStudy (cwrw238):

or you can regard it as (a^3)^3 = a^3 * a^3 * a^3 - adding the exponents this is a^(3+3+3) = a^9

OpenStudy (anonymous):

okay i get it now thank you